pls do this 4me: A TWO DIGIT NUMBERS TENTH PLACE DIGIT IS TWO TIMES GREATER THAN THE UNIT DIGIT. WHEN 8 IS SUBTRACTED FROM THAT NUMBER THE DIGIT BECOMES EQUAL, WHAT IS THAT NUMBER?
Answer by stanbon(75887)

pls do this 4me:
A TWO DIGIT NUMBERS TENTH PLACE DIGIT IS TWO TIMES GREATER THAN THE UNIT DIGIT. WHEN 8 IS SUBTRACTED FROM THAT NUMBER THE DIGIT BECOMES EQUAL, WHAT IS THAT NUMBER?
--------------------
Let the number be 10t + u where t is the tens and u is the units digit.
Equations:
t = 2u
t-8=u
-----
Substitute for "t" and solve for "u":
2u-8 = u
u = 8
t = 2u = 16 (the ten's digit cannot be 16).
--------
I know this is wrong.
What do you mean when you say "When 8 is subtracted from ?that? number
the ?digit? becomes equal? Equal to what?
---
Which number is "that number"?
Which "digit" becomes equal? and equal to what?
-----
Are you posting the problem as it is written in your text?
You should.
Cheers,
Stan H.